solve the following equation in quadratic form: 2x^2/3-3x^1/3-2=0
Answer by lwsshak3(11628)

solve the following equation in quadratic form:
2x^2/3-3x^1/3-2=0
let x^(1/3)=u
x^(2/3)=u^2
..
2u^2-3u-2=0
(2u+1)(u-2)=0
..
2u+1=0
u=-1/2
x^(1/3)=-1/2
cube both sides
x=-1/8
or
u-2=0
u=2
x^(1/3)=2
cube both sides
x=8